Solve the differential equation and find solution
(D2-3DD'+2D'2)=e^2x-y+ex+y
1
Expert's answer
2021-07-30T15:25:13-0400

Hint : When solving the problem, I will assume that on the right side of the equation it looks like this



"\left(D^2-3DD'+2D'^2\right)z=e^{2x-y}+e^{x+y}"

This is an inhomogeneous equation, so its solution consists of two parts



"z(x,y)=(C.F.)+(P.I.)\\[0.3cm]
(C.F.)-\text{Complementary Function of Homogeneous Linear PDE}\\[0.3cm]
(P.I.)-\text{Particular Integral}"

1 STEP : We try to find "(C.F.)"

As we know, the equation



"\boxed{\left(D-mD'\right)z=0\to z(x,y)=\phi(y+mx)}\\[0.3cm]
\phi(y+mx)\,\text{is an arbitrary function}"

In our case,



"\left(D^2-3DD'+2D'^2\right)z=\left(D^2-DD'-2DD'+2D'^2\right)z=\\[0.3cm]
=\left(D\left(D-D'\right)-2D'\left(D-D'\right)\right)z=\\[0.3cm]
=\left(D-D'\right)\left(D-2D'\right)z\\[0.3cm]
\left(D-D'\right)\left(D-2D'\right)z=0\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
\left(D-1D'\right)z=0\to z(x,y)=\phi_1(y+x)\\[0.3cm]
\left(D-2D'\right)z=0\to z(x,y)=\phi_2(y+2x)
\end{array}\right."

Conclusion,



"\boxed{(C.F.)=\phi_1(y+x)+\phi_2(y+2x)}\\[0.3cm]
\phi_1,\phi_2\,\text{are arbitrary functions}"

2 STEP : We try to find "(P.I.)"

As we know



"f\left(D,D'\right)z=F(x,y)\to(P.I.)=\frac{1}{f\left(D,D'\right)}F(x,y)\\[0.3cm]
\text{special case :}\,\,\,\frac{1}{f\left(D,D'\right)}e^{ax+by}=\frac{e^{ax+by}}{f(a,b)},\,\,\,f(a,b)\neq0\\[0.3cm]
if\,\,\,f(a,b)=0 : (P.I.)=x\cdot\frac{1}{f'\left(D,D'\right)}\cdot e^{ax+by},\\[0.3cm]
f'\left(D,D'\right)\text{ is the partial derivative of}\,\,\,f\left(D,D'\right)\,\,\,\text{ w.r.t D}"

In our case,



"\left(D^2-3DD'+2D'^2\right)z=\underbrace{e^{2x-y}}_{(P.I.)_1}+\underbrace{e^{x+y}}_{(P.I.)_2}\\[0.3cm]
(P.I.)_1 : \left(D^2-3DD'+2D'^2\right)z=e^{2x-y}\to\\[0.3cm]
(P.I.)_1=\frac{1}{\left(D^2-3DD'+2D'^2\right)}\cdot e^{2x-y}=\frac{e^{2x-y}}{2^2-3\cdot2\cdot(-1)+2(-1)^2}=\\[0.3cm]
=\frac{e^{2x-y}}{4+6+2}=\frac{e^{2x-y}}{12}\to\boxed{(P.I.)_1=\frac{e^{2x-y}}{12}}\\[0.3cm]
(P.I.)_2=\frac{1}{\left(D^2-3DD'+2D'^2\right)}\cdot e^{x+y}=x\cdot\frac{1}{2D-3D'}\cdot e^{x+y}=\\[0.3cm]
=\frac{x\cdot e^{x+y}}{2-3}=-x\cdot e^{x+y}\to\boxed{(P.I.)_2=-x\cdot e^{x+y}}"

General conclusion.



"\left(D^2-3DD'+2D'^2\right)z=e^{2x-y}+e^{x+y}\to\\[0.3cm]
z(x,y)=(C.F.)+(P.I.)_1+(P.I.)_2\\[0.3cm]
\boxed{z(x,y)=\phi_1(y+x)+\phi_2(y+2x)+\frac{e^{2x-y}}{12}-x\cdot e^{x+y}}\\[0.3cm]
\phi_1,\phi_2\,\text{are arbitrary functions}"

ANSWER



"z(x,y)=\phi_1(y+x)+\phi_2(y+2x)+\frac{e^{2x-y}}{12}-x\cdot e^{x+y}\\[0.3cm]
\phi_1,\phi_2\,\text{are arbitrary functions}"